OpenStudy (anonymous):

Two quantities are related, as shown in the table: x y 2 10 4 9 6 8 8 7 Which equation best represents the relationship?

OpenStudy (campbell_st):

the table is showing a linear relationship. so find the slope by picking any 2 points.... that's the 1st part

OpenStudy (anonymous):

y=1/2x+11 y=-1/2+11 y=2x+10 y=2x+11

OpenStudy (campbell_st):

do you know how to find the slope of a line..?

OpenStudy (anonymous):

i'm trying to figure this out to help show my son how to do I didn't have to learn this when I was in school.

OpenStudy (campbell_st):

ok... so use the points (2, 10) and (4, 9) the slope is the change in the y values divided by the change in x values \[m = \frac{10 - 9}{2 - 4}\] when you calculate the slope you eliminate 2 answers... the the slope intercept form is y = mx + b so substitute the slope you just found then use any point like (6, 8) and substitute these to find b it means x = 6 and y = 8 hope it helps

OpenStudy (anonymous):

ty for explaining it to me.